Calculate the cell potential E°cell of each of the following electrochemical cells at 25C

  1. Sn(s) I Sn2+ (6.5 x 10-3 M) II Ag + (0.110M) I Ag (s)
  2. Zn (s) I Zn 2+ (0.500M) II Fe3+ (7.2 x 10-6 M), Fe2+ (0.15M) I Pt
  3. Pt I H2 (1 atm) I HCl (0.00880M) I Cl2 (1 atm) I Pt
  4. Write the overall cell reaction and calculate the value of E°cell for the following: Zn I Zn2+ II Fe 3+ , Fe2+ I Pt
Answer #1

a)

Sn(s)/Sn+2(6.5 x10^-3M)//Ag+(0.110M)/Ag(s)

according to the given cell representation

oxidation reaction at anode                       Sn(s) ----------------------- Sn+2(aq) + 2e-

reduction reaction at cathode                   [ Ag+(aq) + e- ----------- Ag(s) ] x 2

                                             ----------------------------------------------------------------------------------------

                                                       Sn(s) + 2 Ag+(aq) -------------- Sn+2(aq) + 2 Ag(s)

                                          ---------------------------------------------------------------------------------------------

E0 of Sn+2/Sn = - 0.14V

E0 of Ag+/Ag = + 0.80V

E0cell = E0 cathode - E0anode

E0cell = 0.80 - ( - 0.14)

E0cell = 0.94V

n=2 e-

Ecell = E0cell - 0.0591/n log[Sn+2]/[Ag+]^2

Ecell = 0.94 - 0.0591/2 log( 6.5 x10^-3 )/(0.110)^2

Ecell = 0.948V

Ecell = 0.95V

b)

Zn(s) /Zn+2(0.500M)// Fe+3(7.2x10^-6M),Fe+2(0.15M)/Pt

according to given cell representation

oxidation reaction at anode                        Zn(s) ---------------------- Zn+2(aq)   + 2e-

reduction reaction at cathode                  [ Fe+3(aq) + e- ------------- Fe+2(aq) ] x2

                                          ---------------------------------------------------------------------------------------

                                                  Zn(s) + 2 Fe+3(aq) ---------------- Zn+2(aq) + 2 Fe+2(aq)

                                           ------------------------------------------------------------------------------------------

E0 of Zn+2/Zn = - 0.769V

E0 of Fe+3/Fe+2 = + 0.77V

E0cell = E0cathode - E0 an0de

E0cell = 0.77 - ( - 0.769)

E0cell = 1.539V

n= 2e-

Ecell = E0cell - 0.0591/n log[Zn+2] [ Fe+2]^2/[Fe+3]^2

Ecell = 1.539 - 0.0591/2 log(0.500)(0.15)^2/(7.2x10^-6)^2

Ecell = 1.29V

c) Pt/H2(1 atm)/HCl(0.00880M)/Cl2(1atm)/Pt

according to given cell representation

oxidation at anode                     H2(g) -----------------------2 H+(aq) + 2e-

reduction at cathode                  Cl2(g) + 2e- ----------------- 2 Cl-(aq)

                                           --------------------------------------------------------

                                               H2(g) + Cl2(g) ------------------ 2 H+(aq) + 2Cl-(aq)

                                         -------------------------------------------------------------------

E0 of H+/H2 = 0.00V

E0 of Cl2/Cl- = 1.36

E0cell = E0cathode - E0 anode

E0cell = 1.36 - 0.00

E0cell = 1.36V

Ecell = E0cell - 0.0591/n log[H+]^2[Cl-]^2/[H2][Cl2]

Ecell = 1.36 - 0.0591/2 log(0.00880)^2( 0.00880)^2/1x1

Ecell = 1.60V

d)

Zn(s)/Zn+2//Fe+3,Fe+2/Pt

oxidation reaction at anode                        Zn(s) ---------------------- Zn+2(aq)   + 2e-

reduction reaction at cathode                  [ Fe+3(aq) + e- ------------- Fe+2(aq) ] x2

                                          ---------------------------------------------------------------------------------------

                                                  Zn(s) + 2 Fe+3(aq) ---------------- Zn+2(aq) + 2 Fe+2(aq)

                                           ------------------------------------------------------------------------------------------

E0 of Zn+2/Zn = - 0.769V

E0 of Fe+3/Fe+2 = + 0.77V

E0cell = E0cathode - E0 an0de

E0cell = 0.77 - ( - 0.769)

E0cell = 1.539V